A Bibliography of Books about REDUCE

Anthony C. Hearn The RAND Corporation

Santa Monica, CA 90406-2138 USA

12 December 2023 Version 2.14

Abstract

This bibliography records books and technical reports written about the REDUCE algebra system. ISBN numbers are given, where available.

The bibliography is derived from the books

file available from reduce-netlib@rand.org;

an e-mail message with the text send index

and send index from bibliography will re-

turn indexes of the current holdings.

Title word cross-reference

3.3 [Hea87]. 3.5 [Hea93].

’88 [Mar93].

aided [BC92]. Algebra [DST88, DST93, FBC88, ON90, SHWZ93, BC92, GKW03, HWM93, Ste94, Ueb92, HWM93, MW91, WH88, WH93]. Algebra. [HWM92]. Algebraic [DST88, DST93, Ray87, MW91]. Algorithms [DST88, DST93, SL92, SvW92]. Anwendung [HWM92, HWM93]. Application [WH88, WH93]. applications [GKW03]. Applied [FBC88]. Approach [Mar93].

Beginners [SHWZ93, WH88, WH93]. Brazilian [MW91].

C [Ste94]. Chaos [SvW92]. compact [HWM93]. Complex [Rob15]. Computation [Bee17, DST88, DST93, HI89, NM89, Ray87, SL92]. Computations [SvW92]. Computer [BC92, DST88, DST93, GKW03, MW91, SHWZ93, WH88, WH93, HWM93, Ste94, Ueb92, HWM92, HWM93]. computer-aided [BC92]. Computer-Algebra [HWM93]. Computer-Algebra. [HWM92]. Computeralgebra [Ueb92]. computing [MW91]. course [HWM93].

Design [Mar93]. Doing [HI89]. Dynamics [FBC88, Rob15].

Einführung [Ueb92]. Emergent [Rob15]. energy [Gro97, Gro05]. Evolutionary [Mar93]. Examples [FBC88].

First [MW91]. Fluid [FBC88]. foundations [GKW03]. Fractals [SvW92]. Function [Bee17].

Geophysical [FBC88]. German [HWM93, Ueb92].

Handbook [Bee17, GKW03]. high [Gro97, Gro05].

includes [Ste94]. Introduction [HI89, BC92, Ueb92].

Kompaktkurs [HWM92, HWM93].

lecture [MW91]. Lectures [SHWZ93, WH88, WH93]. Library [Bee17]. Linear [ON90]. LISP [BC92].

Manual [Hea87, Hea93]. MAPLE [Ste94]. MathCW [Bee17]. MATHEMATICA [Ste94]. Mathematical [Bee17]. Mathematical-Function [Bee17]. Mathematics [FBC88, BC92]. Mechanics [NM89, Ste94]. Model [Rob15].

notes [MW91].

PC [HI89]. physics [Gro97, Gro05]. Portable [Bee17]. Program [Mar93]. Programming [Bee17]. programs [Ste94]. pure [BC92].

Quantum [Ste94].

REDUCE [MW91, BC92, FBC88, Gro97, Gro05, Hea87, Hea93, HWM92, HI89, NM89, ON90, Ray87, SL92, Ste94, Ueb92, WH88, WH93, HWM93]. Reuse [Mar93]. RLISP [Mar93].

sample [Ste94]. School [MW91]. Simulation [SHWZ93]. Six [WH88, WH93]. Software [Bee17, Ray87]. Structural [NM89]. Symbolic [FBC88, HI89, NM89]. Systems [DST88, DST93, Rob15, GKW03].

use [HWM93]. User [Hea87, Hea93]. Using [Bee17, Gro97, Gro05, NM89, ON90, Ste94].

Version [Hea87, Hea93].

References

Brackx:1992:CAL

[BC92] F. Brackx and D. Constales. Computer algebra with LISP and REDUCE: an introduction to computer-aided pure mathematics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1441-7. xi + 264 pp. LCCN QA155.7.E4 B72 199.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL

                                                  http://www.springer.com/us/
                                                  book/9783319641096.

Davenport:1988:CAS

[DST88] J. H. Davenport, Y. Siret, and E. Tournier. Computer Algebra, Systems and Algorithms for Algebraic Computation. Academic Press, New York, NY, USA, 1988. ISBN 0-12-204230-1. xix + 267 pp. LCCN QA155.7.E4 D39 1988.

Davenport:1993:CAS

[DST93] J. H. Davenport, Y. Siret, and E. Tournier. Computer Algebra, Systems and Algorithms for Algebraic Computation. Academic Press, New York, NY, USA, second edition, 1993. ISBN 0-12-204230-1. ?? pp. Translated from the French by A. Davenport and J. H. Davenport. Also available in French and Russian.

Flatau:1988:SAA

[FBC88] Piotr J. Flatau, John P. Boyd, and William R. Cotton. Symbolic algebra in applied mathematics and geophysical fluid dynamics —REDUCE examples. Technical Report Version 1.1, Department of Atmospheric Science, Colorado State University, Fort Collins, CO 80521, USA, 1988.

Grabmeier:2003:CAH

[GKW03] Johannes Grabmeier, Erich Kaltofen, and Volker Weispfenning, editors. Computer algebra handbook: foundations, applications, systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-65466-6. xx + 637 pp. LCCN QA155.7.E4 C64954 2003. URL

http://www.springer.com/sgw/
cda/frontpage/0,11855,1-102-
22-1477871-0,00.html. Includes

CD-ROM.

Grozin:1997:URH

[Gro97] A. G. Grozin. Using REDUCE in high energy physics. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-01952-4, 0-521-56002-0 (hardcover). xiv + 384 pp. LCCN QC793.3.H5 G76

1997. URL http://www.inp.nsk.
su/persons/A_G_Grozin/book/.

English translation of an original Russian edition.

Grozin:2005:URH

[Gro05] A. G. Grozin. Using REDUCE in high energy physics. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-01952-4 (paperback). xiii + 384 pp. LCCN QC793.3.H5 G76

                                        2005. URL http://www.loc.gov/
                                        catdir/enhancements/fy0632/
                                        2006272924-d.html;   http://
                                        www.loc.gov/catdir/enhancements/
                                        fy0632/2006272924-t.html.

Hearn:1987:RUM

[Hea87] Anthony C. Hearn. REDUCE User’s Manual, Version 3.3. Technical Report CP 78, Rand Corporation, Santa Monica, CA, USA, July 1987. Also available in a Japanese translation by Hiroshi Toshima from McGraw Hill Japan, 1988.

Hearn:1993:RUM

[Hea93] Anthony C. Hearn. REDUCE User’s Manual, Version 3.5. Technical Report CP 78, Rand Corporation, Santa Monica, CA, USA, October 1993. Also available in a Japanese translation by Hiroshi Toshima from McGraw Hill Japan, 1988.

Hirota:1989:IRD

[HI89] Ryogo Hirota and Masaaki Ito. Introduction to REDUCE — Doing Symbolic Computation on PC. Science sha, Tokyo, Japan, June 1989. In Japanese.

Hehl:1992:CAK

[HWM92] F. W. Hehl, V. Winkelmann, and H. Meyer. Computer-Algebra. Ein Kompaktkurs über die Anwendung von REDUCE. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55724-5. ?? pp.

Hehl:1993:RKA

[HWM93] Friedrich W. Hehl, Volker Winkelmann, and Hartmut Meyer. Reduce: Ein Kompaktkurs über die Anwendung von Computer-Algebra. (German) [Reduce: a compact course on the use of computer algebra]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1993. ISBN 3-642-78227-2, 3-540-56705-4 (print). LCCN QD39.3.E46. URL http://link.

springer.com/openurl?genre=
book\%26isbn=978-3-540-56705-
9;   http://link.springer.com/
10.1007/978-3-642-78227-5.

Marti:1993:REA

[Mar93] Jed B. Marti. RLISP ’88: An Evolutionary Approach to Program Design and Reuse. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1479-0. xiii + 254 pp. LCCN QA76.3.L23 M3 1993.

MacCallum:1991:ACR

[MW91] Malcolm A. H. MacCallum and Francis J. Wright, editors. Algebraic computing with REDUCE: lecture notes from the First Brazilian School on Computer Algebra. Clarendon, New York, NY, USA, 1991. ISBN 0-19-853444-2 (hardcover), 0-19-853443-4 (paperback). LCCN QA155.7.E4 B73 1989.

Nakamura:1989:SCS

[NM89] Hideharu Nakamura and Shouichi Matsui. Symbolic Computation in Structural Mechanics using RE-DUCE. Gihodo Shuppan, Tokyo, Japan, 1989. In Japanese.

Ochiai:1990:LAU

[ON90] Mitsuyuki Ochiai and Kiyokazu Nagatomo. Linear Algebra using REDUCE. Kindai Kagaku sha, January 1990. In Japanese.

Rayna:1987:RSA

[Ray87] Gerhard Rayna. REDUCE— Software for Algebraic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96598-X (New York), 3-540-96598-X (Berlin). ix + 329 pp. LCCN QA155.7.E4 R39 1987.

Roberts:2015:MED

[Rob15] A. J. (Anthony John) Roberts. Model Emergent Dynamics in Complex Systems. Mathematical modeling and computation. SIAM Press, Philadelphia, PA, USA, 2015. ISBN 1-61197-355-4. xii + 748 pp. LCCN QA845 .R62

                                        2015. URL http://www.loc.gov/
                                        catdir/enhancements/fy1503/
                                        2014024843-b.html;   http://
                                        www.loc.gov/catdir/enhancements/
                                        fy1503/2014024843-d.html;

http://www.loc.gov/catdir/enhancements/
fy1503/2014024843-t.html.

Stauffer:1993:CSC

[SHWZ93] D. Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky, editors. Computer Simulation and Computer Algebra. Lectures for Beginners. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56530-2 (New York), 3-540-56530-2 (Berlin). LCCN QA76.9.C65 C656 1993.

Steeb:1992:ACR

[SL92] Willi-Hans Steeb and Dirk Lewien.

Algorithms and Computation with REDUCE. BI-Wissenschaftsverlag, Mannheim, Germany, 1992. ISBN 3-411-15651-1. 153 pp. This book gives a collection of 75 standard methods in mathematics, physics and engineering together with their programs in REDUCE 3.4. Each item consists of a one page mathematical description and one page REDUCE algebraic source text (in most cases a few lines only). The book is an introduction by example to algebraic programming in REDUCE and a collection of ready to use solutions for many mathematical subtasks.

Steeb:1994:QMU

[Ste94] Willi-Hans Steeb. Quantum mechanics using computer algebra: includes sample programs for RE-DUCE, MAPLE, MATHEMAT-ICA and C++. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1770-6. viii + 189 pp. LCCN QC174.17.D37 S74 1994.

Steeb:1992:CFA

[SvW92] Willi-Hans Steeb and Michael Antonie van Wyk. Chaos and Fractals: Algorithms and Computations. BI-Wissenschaftsverlag, Mannheim, Germany, 1992. ISBN 3-411-15961-8. 177 pp. LCCN QA614.86 .S73 1992. This book includes 26 Turbo C and C++, 47 Turbo PASCAL, and 23 REDUCE programs.

Ueberberg:1992:ECR

[Ueb92] Johannes Ueberberg. Einführung in die Computeralgebra mit RE-DUCE. (German) [Introduction to computer algebra with RE-DUCE]. BI-Wissenschaftsverlag, Mannheim, Germany, 1992. ISBN 3-411-15781-X. 331 pp.

Winkelmann:1988:RBS

[WH88] Volker Winkelmann and Friedrich W. Hehl. REDUCE for Beginners, Six Lectures on the Application of Computer Algebra. In D. Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky, editors, Computer Simulation and Computer Algebra. Lectures for Beginners. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96598-X (New York), 3-540-96598-X (Berlin).

Winkelmann:1993:RBS

[WH93] Volker Winkelmann and Friedrich W. Hehl. REDUCE for Beginners, Six Lectures on the Application of Computer Algebra. In Stauffer et al. [SHWZ93], page ?? ISBN 0-387-56530-2 (New York), 3-540-56530-2 (Berlin). LCCN QA76.9.C65 C656 1993.